Find five pairs of numbers (a,b) which are solutions to ab = 70. (Note that you can use negative numbers and fractions
Answer by rothauserc(4718)

the prime factors of 70 are
2, 5, 7
therefore we have
14*5 = 70
2*35 = 70
10*7 = 70
70*1 = 70
-70*-1 = 70